OpenStudy (anonymous):

@campbell_st can you help me with this one as well. The length of a rectangle is 4 cm less than the second less than three times the width. The perimeter is 64 cm. Find the width, length and equation. (Hint: Perimeter=21+2w)

OpenStudy (campbell_st):

sorry the dimensions don't make a lot of sense. but start by letting the width = x and then find the length... then apply the formula.

OpenStudy (anonymous):

length= 3w-4

OpenStudy (anonymous):

the perimeter is 64 cm

OpenStudy (campbell_st):

ok... so then 64 = 2w + 2(3w -4) solve for w, but distribute 1st

OpenStudy (anonymous):

w= 9

OpenStudy (campbell_st):

yes, that makes sense 64 = 2(9) + 2(27 - 4) 64 = 18 + 46

OpenStudy (anonymous):

thank you :)